2. Sulfur dioxide gas and oxygen gas react to produce gaseous sulfur trioxide. What mass  of sulfur trioxide is produced when 125 g of sulfur dioxide reacts with 50.00 g of  oxygen to produce sulfur trioxide as the only product


1
Expert's answer
2021-01-18T01:52:36-0500

SO2 + 1/2 O2 = SO3

n(SO2) = (125 g)/(64.066 g/mol) = 1.9511 mol

n(O2) = (50 g)/(32 g/mol) = 1.5625 mol

SO2 is the limiting reactant

n(SO3) = n(SO2) = 1.9511 mol

m(SO3) = (1.9511 mol)*(80.06 g/mol) = 156.2 g